A cyclist traveled at a rate of 24 mph to visit a nearby town. The
cyclist averaged 18 mph on the return trip. if the round trip took 4.9
hours, find the distance to the nearby town.

Answer by checkley71(8403)

DISTANCE=RATE*TIME OR TIME=DISTANCE/RATE SO THE TIME FOR THE 2 TRIPS
D/24+D/18=4.9 HOURS OR (18D+24D)/24*18=4.9 OR 42D/432=4.9 OR 42D=432*4.9 OR
42D=2116.8 OR D=2116.8/42 OR D=50.4 MILES TO THE NEARBY TOWN